# Near-field Beamforming for Extremely Large-scale MIMO Based on Unsupervised Deep Learning

### v1

@@ II-C Near-field Beamforming
We presume the utilization of a predefined codebook for beamforming. Considering that the steering vector in the near-field is influenced by both angle and distance, the angular domain Discrete Fourier Transform (DFT) codebook utilized in the far-field domain becomes obsolete. Upon incorporating distance sampling, the codebook in the polar coordinate domain is expressed as:
$\mathcal{W}=$$\left\{\mathbf{b}\left(\theta_{1},r_{1}\right),\ldots,\mathbf{b}\left(\theta_{1},r_{j}\right),\ldots,\mathbf{b}\left(\theta_{1},r_{J}\right),\right.$ (11)
@@
$\left.\mathbf{b}\left(\theta_{i},r_{1}\right),\ldots,\mathbf{b}\left(\theta_{i},r_{j}\right),\ldots,\mathbf{b}\left(\theta_{i},r_{J}\right),\right.$
@@
$\left.\mathbf{b}\left(\theta_{I},r_{1}\right),\ldots,\mathbf{b}\left(\theta_{I},r_{j}\right),\ldots,\mathbf{b}\left(\theta_{I},r_{J}\right)\right\},$
where each element represents a candidate beam code word, $\theta_{i}$ denotes the $i$-th sampling angle, and $r_{j}$ denotes the $j$-th sampling distance. Our objective is to optimize the selection of codewords from a predefined codebook to maximize the achievable data rate. The pertinent problem of code word selection can be formulated as:
$\boldsymbol{w}^{*}=\text{arg}\max_{\boldsymbol{w}\in\mathcal{W}}\log_{2}\left(1+\frac{P\left|\boldsymbol{hw}\right|^{2}}{\sum_{i\neq m}{\left|\boldsymbol{w}^{H}\boldsymbol{h}_{i}\right|}^{2}+\sigma_{n}^{2}}\right).$ (12)
The most straightforward approach to address the aforementioned challenges is through exhaustive beam scanning, aiming to find the optimal beam configuration that maximizes the achievable data transfer rate. However, compared to far-field scenarios, this method may incur unacceptable training costs, especially in mmWave communications, where higher frequencies and increased signal blockage lead to elevated path losses. Moreover, the nonlinear distance information in the near-field channel further complicates beam training, making traditional methods challenging for accurate modeling.
Given the outstanding performance of deep learning in handling complex nonlinear problems, particularly when trained on large-scale datasets, it has proven effective in uncovering intricate patterns hidden within the data. In the next section, we will explore how to leverage deep learning techniques, specifically neural networks, to accomplish beam training in near-field channels. Through deep learning, our aim is to more effectively capture the nonlinear characteristics of the channel, thereby reducing the computational and communication overhead associated with beam training.
III Deep learning based near-field beamforming Method
III-A Problem Formulation
The near-field codebook typically experiences a significant size due to the close correlation between the near-field channel and both angle and distance information. This introduces substantial computational complexity when employing codebook-based beam design. In practical scenarios, our primary focus is on the beam gain directed towards the target UE’s location, rather than optimizing the beam gain for every grid across the entire space. Our objective is to achieve effective control over the beamforming shape without relying on a predefined codebook, aiming to maximize the achievable data transfer rate. Considering the constant modulus constraint $\lVert w_{i}\rVert^{2}=1,\ \text{for}\,\,i=1,\dotsi,N_{t}$, the optimization problem concerning the beamforming vector $\boldsymbol{w}$ is given by the following formula:
@@ III-A Problem Formulation
s.t. $w_{i}\,\,\in\boldsymbol{w}$
$\left|{{w}}_{i}\right|^{2}=1,\forall\ {{i}\in N_{t}}.$
The channel state information in XL-MIMO systems is intricate, giving rise to the nonlinearity and significant non-convexity of the aforementioned challenges.
We introduce deep learning as a powerful tool to address these issues. In this paper, a unsupervised learning of neural networks model is designed to enable beamforming vectors to adapt and learn flexibly under complex channel conditions. Our goal is to find the mapping function $\mathbf{\Phi_{\xi}}$ of parameter $\boldsymbol{\xi}$, which uses CSI to predict the best beamforming vectors $\boldsymbol{w}$. The mapping function can be expressed as:
@@
where $\boldsymbol{h}$ denotes the downlink channel state information from the BS to the UE, and $\boldsymbol{\hat{w}}$ represents the desired beamforming vector. It should be noted that the process of obtaining $\boldsymbol{\hat{w}}$ from $\boldsymbol{h}$ is quite complicated.
III-B Deep Learning Scheme Design
@@
The overall architecture of our proposed model is shown in Fig. 4. The channel sample $h$ generated by the simulation is input into the neural network. Subsequently, the neural network iteratively learns the model parameters through gradient descent to minimize the predefined loss function. By adjusting the learning rate, we can converge the model to an approximate global optimal solution.
Once the model is trained, all parameters of the network $\xi$ will be fixed. In online deployment, we only need to input the channel status information $\boldsymbol{h}$, and the network will directly output the beamforming vector $\boldsymbol{w}$. Algorithm 1 shows the pseudo-code of our proposed scheme. In the design of the network structure, we introduce some innovative mechanisms to ensure the effectiveness of the designed scheme.
@@
1 Initialize: The number of UEs $K$; SNR; The number of transmit antennas $N_{t}$; The normalized transmit power $P$;
2 Calculate the input $\boldsymbol{h}$ of neural network according to formula 7. Train the Neural Network:;
@@
4      Randomly sample a batch of training data $(h,\text{SNR})$;
@@ 6       Calculate the phase $\boldsymbol{\theta}\leftarrow\pi\boldsymbol{\theta}$;
7       Calculate the complex beamforming vector according to formula 16;
8       Calculate the loss according to formula 17;
@@ 13 Given new channel state information $\boldsymbol{h_{\text{new}}}$;
Algorithm 1 Unsupervised Deep Learning for Near-field Beamforming
1) Input Data Generation: Each simulation generates positions for K UEs at T time frames, all of which fall within the specified Angle and distance sampling range. When the BS provides service to UE $k$, the remaining K-1 UEs are considered interference, which will be reflected in the loss function we design. The channel sample h generated by the system model presented in Section 3 is used as input data. Detailed parameter design is shown in Table II.
@@
Number of Transmit Antennas $N_{t}$ 256
Number of Receive Antennas $N_{r}$ 1
@@
Sampling Angle $\theta$ [-60°, 60°]
@@
Figure 5: Illustrate the structure of the convolutional neural network design. The red dotted boxes depict the intricate details of feature extraction for the aforementioned complex data. The orange rectangle represents the data transformations within the convolutional neural network, and the red rectangle highlights the process of obtaining the desired phase $\theta$.
@@
2) Complex Number Feature Extraction: Different from previous approaches that solely rely on multi-layer fully connected methods, we leverage convolutional neural networks for the intricate task of extracting features from complex signals. The input data dimension is $2\times N_{t}$, and we transform it into $1\times 2\times N_{t}$ by adding an extra dimension. The structure is shown in Fig. 5, the orange rectangles depict the layers of our model. Each set of three identically sized rectangular blocks is termed a feature extraction block. Within this block, the data undergoes two convolution layers, a regularization layer (BatchNorm), and an activation layer (ReLU).
@@
With meticulously designed padding and convolution kernel size, we ensure that the data size remains unchanged before and after passing through the feature extraction block, as outlined in the red dashed box. Initially, the input data of $B\times 1\times 2\times N_{t}$ is transformed to $B\times 1\times 4\times(N_{t}+2)$ by introducing padding, followed by a 2×2 convolution operation. In the convolution process, the red position of the kernel extracts real part information, the blue position captures information combining real and imaginary parts, and the green position extracts imaginary part information. After the initial convolution operation, the data size becomes $B\times 1\times 3\times(N_{t}+1)$. Subsequent convolution restores the data size to $B\times 1\times 2\times N_{t}$, aligning with the dimensions of the input data.
@@
After the data undergoes a feature extraction block, we employ average pooling for downsampling. At this point, the channel count increases, and the data length is halved compared to the original. After repeating this process twice, the data size becomes $B\times c\times 2\times N_{t}/4$, where c is the number of channels. Subsequently, the data undergoes deconvolution and a feature extraction block, where the channel count is reduced, and the data length is doubled. After repeating this procedure twice, the data size becomes $B\times 1\times 2\times N_{t}$, matching the dimensions of the input data. The utilization of a convolutional neural network not only allows for the extraction of more diverse data features but also significantly reduces the model’s complexity. The analysis of model complexity is presented in the following section.
@@
3) Convert Angles into Beamforming Vectors: In the left end of Fig. 5, after the feature extraction block fully captures the features, the data is flattened into a one-dimensional vector. The data size becomes $1\times N_{t}$ following a fully connected layer. To ensure that the neural network’s output is a complex-valued vector adhering to the norm constraint, a ”tanh” layer is introduced at the final layer of the network. Specifically, its output is a real value denoted as $\theta$, and after passing through the ”tanh” activation layer, its value is confined to the range $(0,1)$. Multiplying the output by $\pi$ yields the phase within the range of $[-\pi,\pi]$. The mathematical expression for this process is as follows:
@@
Then, Euler’s formula is applied to convert the phase into the desired complex-valued beamforming vector, with the complex output given by the following formula:
$w=\exp\left(\text{j}\cdot\pi\right)=\cos\left(\pi\boldsymbol{\theta}\right)+\text{j}\cdot\sin\left(\pi\boldsymbol{\theta}\right),$ (16)
where $j=\sqrt{-1}$. It can be observed that $\theta\pi$ has a clear physical meaning, with its elements corresponding to the equivalent phases of the beamforming coefficients in $w$.
4) Loss Function Design: We adopt an unsupervised learning approach, where the neural network is guided by a loss function directly linked to the beamforming target. The loss function for the task is defined as the negative of the average achievable rate.
@@
$=-\frac{1}{QK}\sum_{q=1}^{Q}\sum_{k=1}^{K}\log_{2}\left(1+\frac{\left|w^{H}h_{k}\right|^{2}}{\sum_{i\neq m}\left|w^{H}h_{i}\right|^{2}+\sigma_{n}^{2}}\right)$
Where $Q$ represents the total number of training samples and $k$represents the number of UEs. $\sigma_{n}^{2}$, $h_{k}$, and $w$ represent the SNR, CSI, and the output analog beamforming vector associated with the Kth sample, respectively. It should be emphasized that the decrease in loss directly corresponds to the increase in the average achievable rate.
IV Experimental Setup and Simulation Analysis
IV-A Experimental Setup
@@
In the proposed unsupervised deep learning model, the hyperparameter Settings are shown in Table III. The data set is randomly divided into a 75% training set and a 25% test set. The learning rate is initialized to $10^{-2}$ and dynamically decreases when the evaluation metric no longer improved. Parameter tuning is conducted using the Adam optimizer, with a batch size set to $256$.
@@
IV-B Experimental Result
During the simulation experiment, a evenly spaced half-wave linear array with Nt = 256 was deployed at BS. The simulation parameters align with the millimeter-wave near-field channel model outlined in Section 3. The results of the proposed scheme compared with the baseline scheme [35, 36] are as follows, with the average achievable rate defined as the evaluation indicator.
@@ IV-B Experimental Result
The model training loss curve is presented in Fig. 6 to validate its effective convergence. The four curves depict the loss evolution of the training set and the validation set with the tanh layer, and the loss changes of the training set and the validation set without the tanh layer, respectively. It can be observed that models with tanh layers exhibit superior performance, with lower loss values, and also better performance on verification sets. Conversely, models without tanh layers exhibit pronounced overfitting. This confirms that the incorporation of the tanh layer to confine the phase within the range [-1, 1] enhances the nonlinearity of the model and effectively mitigates overfitting.
@@
We rigorously examine whether the inclusion of the tanh layer can enhance the beamforming performance. Fig. 7 illustrates the average achievable rate associated with the beamforming vector output by the model before and after the addition of the tanh layer. The results show that for all SNR values, the model with tanh layer achieves a higher average achievable rate than the model without tanh layer. And this improvement becomes more pronounced with increasing SNR.
Fig. 8 shows the variation in the average achievable rate with the SNR under different beam training schemes, where SNR increases from -20dB to 20dB. The graph clearly demonstrates that the proposed scheme consistently outperforms existing far-field and near-field beam training schemes across a wide range of SNR values. When SNR = 20dB, our proposed scheme achieves an approximately 18%, 30%, and 120% improvement in average achievable rate compared to the near-field hierarchical beamforming scheme, far-field hierarchical scheme, and exhaustive search scheme, respectively. Notably, when the SNR$>$-5dB, the scheme shows obvious performance gain compared with the comparison scheme. Furthermore, the proposed scheme does not exhibit clear superiority at low SNR. Under low SNR conditions, where noise power surpasses signal power, neural networks struggle to effectively extract signal features. Our scheme might be vulnerable to noise under conditions of low SNR. Investigating potential solutions to mitigate noise sensitivity represents a focal point for our future research efforts.
@@
With the expanding communication frequency band, carrier frequency has emerged as a crucial factor influencing communication rates. Illustrated in Fig. 9 are the average achievable rate curves for the proposed scheme and the comparative scheme at different central carrier frequencies. In this context, the SNR is set to 10dB, and UE’s positions and angles align with the specifications outlined in Table II.The figure distinctly illustrates that for all carrier frequencies, our proposed scheme consistently outperforms other schemes in terms of average achievable rates. For the far-field hierarchical scheme, with the increase of carrier frequency, near-field propagation gradually dominates, resulting in that the average achievable rate of the far-field scheme basically no longer increases when the carrier is greater than 40GHz. Optimal performance for all solutions is observed at a carrier frequency of 50GHz. This is attributed to the increased free-space propagation losses incurred by high-frequency signals during transmission. As the carrier frequency rises, the performance gain achieved by augmenting the number of antennas becomes insufficient to counterbalance the escalating free-space propagation losses. Notably, when $f_{c}>$ 50GHz, the comparative scheme experiences a decline in performance, whereas the proposed scheme’s performance remains relatively stable, affirming the robustness of our proposed solution. Given the constraint of restricting the training cost for the exhaustive search scheme to 256, the scheme encounters challenges in conducting precise searches in high-frequency bands, leading to a notable degradation in performance.
Figure 10: Average achievable rate curves at different distance.
@@
Fig. 10 illustrates the correlation between the average achievable rate and the UE-to-BS distance. Here, the distance r from the UE to the BS gradually increases from 5 meters to 50 meters. The SNR is set to 10dB, the UE’s angle $\theta$ is set [-60°, 60°], and the carrier frequency is set to 50GHz. Other simulation parameters are consistent with those in Table II. The observations in Fig. 10 reveal the substantial impact of near-field propagation on beam training performance. Notably, when the UE is in close proximity to the BS, our proposed scheme exhibits superior rate performance compared to all the schemes under consideration. However, as the distance between the UE and the BS increases, the performance degradation of our scheme becomes comparable to that of the far-field hierarchical beam training scheme. For far-field schemes, the dominance of near-field propagation as the distance decreases leads to a rapid decline in the average rate of far-field schemes. Due to the limitation of the maximum training cost, the exhaustive scheme experiences severe degradation in the near-field region, impeding effective search for near-field positions. Additionally, the near-field hierarchical beam training scheme can moderately alleviate the decline in the average achievable rate, but its performance gradually diminishes when the distance exceeds 30 meters. Furthermore, when the distance is less than 20 meters, its performance notably lags behind the proposed scheme. In contrast, the proposed scheme demonstrates the ability to search for the optimal beamforming vector with minimal pilot overhead. This affirms the robustness of the proposed scheme in both near and far-field communications scenarios.
@@
Fig.11 depicts the variation in average achievable rate with angle $\theta$. Here, $\theta$ gradually increases from -60 °to 60°, and the distance ($r$) is randomly generated within the range [5m, 50m]. The SNR is set to 10 dB, and the remaining simulation settings mirror those in Table II. Notably, for the far-field scheme, performance degrades significantly as $\theta$ approaches 0. This degradation is attributed to the increased prominence of near-field effects near zero angles, which has been demonstrated in the literature [37]. Furthermore, the near-field hierarchical beam training scheme exhibits noticeable fluctuations. This is because the near-field mixing scheme creates a near-field codebook by uniformly sampling angles and distances in Cartesian coordinates. This approach proves challenging in achieving stable beamforming performance across the entire near-field environment [30]. In contrast, our proposed scheme consistently attains the highest average achievable rate across all considered angles. his result underscores the robustness and effectiveness of our approach in handling the challenges posed by varying angles in near-field communications scenarios.

### v2

@@ III-B Near-field Beamforming
We presume the utilization of a predefined codebook for beamforming. The codebook in the polar coordinate domain is expressed as:
$\mathcal{W}=$$\left\{\mathbf{b}\left(\alpha_{1},r_{1}\right),\ldots,\mathbf{b}\left(\alpha_{1},r_{j}\right),\ldots,\mathbf{b}\left(\alpha_{1},r_{J}\right),\right.$ (8)
@@
$\left.\mathbf{b}\left(\alpha_{i},r_{1}\right),\ldots,\mathbf{b}\left(\alpha_{i},r_{j}\right),\ldots,\mathbf{b}\left(\alpha_{i},r_{J}\right),\right.$
@@
$\left.\mathbf{b}\left(\alpha_{I},r_{1}\right),\ldots,\mathbf{b}\left(\alpha_{I},r_{j}\right),\ldots,\mathbf{b}\left(\alpha_{I},r_{J}\right)\right\},$
where each element represents a candidate beam code word, $\alpha_{i}$ denotes the $i$-th sampling angle, and $r_{j}$ denotes the $j$-th sampling distance. Generally, the optimal codeword is selected from a predefined codebook to maximize the achievable data rate. Problems related to code word selection can be expressed as:
$\boldsymbol{w}^{*}=\text{arg}\max_{\boldsymbol{w}\in\mathcal{W}}\log_{2}\left(1+\frac{G\left|\boldsymbol{w}^{H}\boldsymbol{h}_{k}\right|^{2}}{\sum_{i\neq k}{\left|\boldsymbol{w}^{H}\boldsymbol{h}_{i}\right|}^{2}+\sigma_{n}^{2}}\right).$ (9)
The most straightforward approach is to find the optimal beam through exhaustive beam scanning. However, this method may incur unacceptable training costs compared to far-field scenarios. Moreover, the nonlinear distance information in the near-field channel further complicates beam training, making traditional methods challenging for accurate modeling.
Deep learning excels at dealing with complex nonlinear problems. In the next section, we will explore how to leverage deep learning to accomplish beam training in near-field channels. Our aim is to more effectively capture the channel’s nonlinear characteristics and reduce the computational and communication overhead associated with beam training.
IV Deep learning based near-field beam training Method
IV-A Problem Formulation
IV-B Problem Formulation
Near-field beamforming usually has a large codebook size, which means using a codebook-based beam training scheme will incur huge training overhead. In practical scenarios, our primary focus is on the beam gain directed toward the target UE’s location. We aim to achieve effective beamforming without relying on a predefined codebook to maximize the achievable data transfer rate. Considering the constant modulus constraint $\lVert w_{i}\rVert^{2}=1,\ \text{for}\,\,i=1,\dotsi,N$, the optimization problem $\boldsymbol{w}$ is given by the following formula:
@@ IV-B Problem Formulation
s.t. $\lVert w_{i}\rVert^{2}=1,\forall\ {{i}\in\mathcal{N}=\{1,2,\ldots,N\}}.$
The CSI in XL-MIMO systems is intricate, giving rise to the nonlinearity and significant non-convexity of the aforementioned challenges.
We introduce deep learning as a powerful tool to address these issues. In this paper, a deep neural network model is designed to flexibly adapt and learn beamforming vectors under complex channel conditions. Our goal is to find the mapping function $\mathbf{\Phi_{\xi}}$ taking $\boldsymbol{\xi}$ as the parameter, which uses CSI to predict the optimal beamforming vector $\boldsymbol{w}$. The mapping function can be expressed as:
@@
where $\boldsymbol{h}$ denotes the downlink CSI from the BS to the UE, and $\boldsymbol{\hat{w}}$ represents the desired beamforming vector.
IV-C Deep Learning Scheme Design
@@
1 Initialize: The number of UEs $K$; SNR; The number of transmit antennas $N$; The normalized transmit power $P$;
2 Calculate the input $\boldsymbol{h}$ of neural network according to formula 4. Train the Neural Network:;
@@
4      Randomly sample a batch of training data $(\boldsymbol{h},\text{SNR})$;
@@ 6       Calculate the phase $\boldsymbol{\theta}\leftarrow\pi\boldsymbol{\theta}$;
7       Calculate the complex beamforming vector according to formula 13;
8       Calculate the loss according to formula 14;
@@ 13 Given new channel state information $\boldsymbol{h_{\text{new}}}$;
Algorithm 1 Near-field Beam Training Based on Deep Learning
The overall architecture of our proposed model is shown in Fig. 4. The channel sample $\boldsymbol{h}$ is input into the neural network. Subsequently, the neural network iteratively learns the model parameters through gradient descent to minimize the predefined loss function. We can converge the model to an approximate global optimal solution by adjusting the learning rate.
Once the model is trained, all parameters of the network $\xi$ will be fixed. In online deployment, we only need to input the channel sample $\boldsymbol{h}$, and the network will directly output the beamforming vector $\boldsymbol{w}$. Algorithm 1 shows the pseudo-code of our proposed model. In the design of the network structure, we introduce some innovative mechanisms to ensure the effectiveness of the designed model.
1) Input Data Generation: Each simulation generates positions for $K$ UEs at $T$ time frames, all of which fall within the specified angle and distance sampling range. When the BS provides service to UE $k$, the remaining $K-1$ UEs are considered as interference. The channel sample $h$ presented in Section IV is used as input data. Detailed parameter design is shown in Table II.
@@
Number of Transmit Antennas $N$ $256$
@@
Sampling Angle $\alpha$ [$-60$°, $60$°]
@@
Figure 5: Illustrate the structure of the convolutional neural network design. The red dashed box depict the details of the feature extraction block. The orange rectangle represents the data transformations in the convolutional neural network.
@@
2) Complex Number Feature Extraction: Different from previous approaches that solely rely on fully-connected layers, we leverage convolutional neural networks to extract features from complex signals. The input data dimension is $2\times N$, and we transform it into $1\times 2\times N$ by adding an extra dimension. Fig. 5 shows the structure of designed convolutional neural networks. A feature extraction block contains two convolution layers, two batch normalization layers (BN), and two rectified linear unit (ReLU).
@@
By strategically designing the padding and convolution kernel size, we ensure that the data size remains constant before and after passing through a feature extraction block, as shown in the red dashed box in Fig. 5. Initially, convert the input data of size $B\times 1\times 2\times N$ to $B\times 1\times 4\times(N+2)$ by adding padding. Then, $2\times 2$ convolution layers are used for feature extraction. In the convolution process, the red convolution kernel extracts real part information, the green convolution kernel extracts imaginary part information, and the blue convolution kernel extracts the information combining the real and imaginary parts (phase). After one convolutional layer, the data size becomes $B\times 1\times 3\times(N+1)$. The subsequent convolutional layer restores the data size to $B\times 1\times 2\times N$, aligning with the dimensions of the input data.
@@
After the data undergoes a feature extraction block, we employ average pooling for downsampling. At this point, the channel count increases, and the data length is halved compared to the original. After repeating this process twice, the data size becomes $B\times C\times 2\times N/4$, where $C$ is the number of channels. Subsequently, after two deconvolution and two feature extraction blocks, the data size becomes $B\times 1\times 2\times N$, matching the dimension of the input data.
@@
3) Convert Angles into Beamforming Vectors: In the left end of Fig. 5, the data is flattened into a one-dimensional vector after the feature extraction block fully captures the features. The data size becomes $1\times N$ following a fully connected layer. To ensure that the neural network’s output is a complex-valued vector adhering to the norm constraint, the “tanh” layer is utilized at the final layer of the network. Specifically, its output is a real value denoted as $\theta$. After passing through the “tanh” activation layer, its value is confined to the range $(-1,1)$. Multiplying the output by $\pi$ yields the phase within the range of $[-\pi,\pi]$. The mathematical expression for this process is as follows:
@@
Then, Euler’s formula is applied to convert the phase into the desired complex-valued beamforming vector:
$\boldsymbol{w}=\exp\left(\text{j}\cdot\pi\boldsymbol{\theta}\right)=\cos\left(\pi\boldsymbol{\theta}\right)+\text{j}\cdot\sin\left(\pi\boldsymbol{\theta}\right),$ (13)
where $j=\sqrt{-1}$. It can be observed that $\theta\pi$ has a clear physical meaning, with its elements corresponding to the equivalent phases of $w$.
4) Loss Function Design: We adopt a codebook-free learning approach, where the training of the neural network is guided by a loss function directly linked to the beamforming target. The loss function for the task is defined as the negative of the average achievable rate.
@@
$=-\frac{1}{Q}\sum_{q=1}^{Q}\log_{2}\left(1+\frac{\left|w^{H}h_{k}\right|^{2}}{\sum_{i\neq k}\left|w^{H}h_{i}\right|^{2}+\sigma_{n}^{2}}\right)$
Where $Q$ represents the total number of training samples and $k$ represents the number of UEs. $\sigma_{n}^{2}$, $h_{k}$, and $w$ represent the SNR, CSI, and the output analog beamforming vector associated with the $K$th sample, respectively. Notably, the decrease in loss directly corresponds to the increase in the average achievable rate.
V Experimental Setup and Simulation Analysis
V-A Experimental Setup
@@
The hyperparameter settings in the proposed deep learning model are shown in Table III. The data set is randomly divided into a $75\%$ training set and a $25\%$ test set. The batch size is set to $256$. The learning rate is initialized to $10^{-2}$ and dynamically decreases when the evaluation metric is no longer improved. Parameter tuning is conducted using the Adam optimizer.
@@
V-B Experimental Result
An evenly spaced half-wave linear array with $Nt=256$ was deployed at BS during the simulation experiment. The simulation parameters align with the millimeter-wave near-field channel model outlined in Section IV. The average achievable rate is defined as the evaluation indicator. The compared benchmarks are as follows:
  • Near-field hierarchical beam training: The hierarchical near-field codebook comprises several levels of sub-codebooks, which are determined by different sampling angles and distances. We sequentially search from the first-level sub-codebook to the last, where the sampling rate of each subsequent level is influenced by the optimal codeword found in the previous level [15]. Throughout the beam training process, the sampling angles and distances gradually narrow down, ultimately yielding the global optimal codeword.
  • Far-field hierarchical beam training: This is a classical beam training method for far-field scenarios. At the training stage, different levels of codebooks are used for beam alignment, progressively narrowing the beam scanning angle to find the optimal beam direction [39].
  • Exhaustive search beam training: This scheme simultaneously samples multiple angles $\alpha\in[\alpha_{min},\alpha_{max}]$ and distances $r\in[\rho_{min},+\infty]$ to construct the near-field codebook. It first fixes the distance loop and exhaustively searches all angles. Then, change the distance loop and repeat the above steps until the entire codebook is traversed to obtain the optimal beam focusing vector [36]. Exhaustive search beam training incurs significant training overhead, which can be mitigated by reducing the sampling rate but at the cost of reduced beamforming performance. We simulate the exhaustive search beam training schemes with a training cost limit of $256$ and unlimited, respectively.
@@ V-B Experimental Result
The model training loss curve is presented in Fig. 6 to validate its effective convergence. These four curves describe the loss evolution of the training and validation sets with and without tanh, respectively. It can be observed that models with tanh layers exhibit superior performance, with lower loss values, and also better performance on verification sets. Conversely, models without tanh layers exhibit pronounced overfitting. This confirms that the incorporation of the tanh layer to confine the phase within the range [-1, 1] enhances the nonlinearity of the model and effectively mitigates overfitting.
@@
We examine whether the inclusion of the tanh layer can enhance the beamforming performance. Fig. 7 shows the average achievable rate associated with the beamforming vector output by the model before and after the addition of the tanh layer. The results show that for all SNR values, the model with tanh layer achieves a higher average achievable rate than the model without tanh layer. And this improvement becomes more pronounced with increasing SNR.
Fig. 8 shows the variation in the average achievable rate with the SNR under different beam training schemes. It demonstrates that the proposed scheme performs close to the unlimited exhaustive search scheme and consistently outperforms existing far-field and near-field hierarchical beam training schemes across a wide range of SNR values. When SNR = $20$dB, our proposed scheme achieves an approximately $18\%$, $30\%$, and $120\%$ improvement in average achievable rate compared to the near-field hierarchical beam training scheme, far-field hierarchical scheme, and $256$ overhead exhaustive search scheme, respectively. Notably, when the SNR$>-5$dB, the scheme shows an obvious performance gain compared with the comparison scheme. The observed performance enhancement is mainly due to the model aiming to maximize the average achievable rate. The scheme design of codebook-free allows for more dynamic and efficient beamforming compared to traditional methods. Furthermore, the proposed scheme does not exhibit clear superiority at low SNR. Under low SNR conditions, where noise power surpasses signal power, neural networks struggle to effectively extract signal features. Our scheme might be vulnerable to noise under conditions of low SNR. Investigating potential solutions to mitigate noise sensitivity represents a focal point for our future research efforts.
@@
Carrier frequency is a crucial factor influencing communication rate. Fig. 9 shows the average achievable rate curves of the proposed scheme and the compared scheme at different carrier frequencies. The figure illustrates that for all carrier frequencies, the average achievable rate of our proposed scheme consistently outperforms other schemes. With the increase of carrier frequency, near-field propagation gradually dominates, resulting in the average achievable rate of the far-field scheme no longer increasing when the carrier is greater than $40$ GHz. Optimal performance for all solutions is observed at the carrier frequency of $50$ GHz. This is attributed to the increased FSPL of high-frequency signals during transmission. As the carrier frequency increases, the gain from increasing the number of antennas will not be enough to offset FSPL. Notably, when $f_{c}>$ $50$ GHz, the performance of the proposed scheme performance remains stable, affirming the robustness of our proposed solution. Given the overhead of the exhaustive search scheme is limited to $256$, the scheme encounters challenges in conducting precise searches in high-frequency bands, leading to a notable degradation in performance.
Figure 10: (a) and (b) represent the variation of the average reachable rate with user distance without and with path loss, respectively.
@@
Fig. 10 illustrates the average achievable rate against the UE-to-BS distance. Here, the distance $r$ from the UE to the BS gradually increases from $5$ to $50$ meters, and the SNR is set to $10$ dB. As can be seen from Fig. 10 (a), when FSPL is not considered, the average achievable rate of our proposed scheme is superior to all the compared schemes in the near-field region. As the distance increases, the performance of our proposed scheme is comparable to that of the far-field hierarchical beam training scheme. Since only angle search can be performed, the performance of far-field hierarchical schemes deteriorates rapidly in the near-field region. Due to the maximum training cost limitation, the $256$ exhaustive scheme also experiences severe degradation in the near-field region, impeding the practical search for near-field positions. Additionally, the near-field hierarchical beam training scheme can moderately alleviate the decline in the average achievable rate, but its performance gradually diminishes when the distance exceeds $30$ meters. Fig. 10 (b) shows that due to FSPL, the performance of all beam training schemes decreases as the distance increases. However, the proposed beam training scheme still maintains high performance at all distances and is close to the performance of the exhaustive scheme without limited training cost. In contrast, our proposed scheme demonstrates the ability to search for the optimal beamforming vector with minimal pilot overhead. This affirms the robustness of the proposed scheme in both near and far-field communications scenarios.
@@
Fig.11 depicts the variation in average achievable rate with angle $\alpha$. Here, $\alpha$ gradually increases from $-60$°to $60$°, and the SNR is set to $10$ dB. Notably, for the far-field scheme, performance degrades significantly as $\alpha$ approaches 00. This degradation is attributed to the increased prominence of near-field effects near zero angles, as demonstrated in the literature [40]. Furthermore, the near-field hierarchical beam training scheme exhibits noticeable fluctuations. This is because the near-field hierarchical scheme creates the codebook by uniformly sampling angles and distances in cartesian coordinates. This approach proves challenging in achieving stable beamforming performance across the entire near-field environment [34]. In contrast, our proposed scheme consistently attains the high average achievable rate across all angles. This underscores the robustness and effectiveness of our approach in handling the challenges posed by near-field communications scenarios.